Solve (a) 〖 4〗^x=8,

(b) 〖 2〗^8÷8=2^(x-1)

Answer by stanbon(75887)

I'm going to assume your symbol means "absolute value".
-----------------------
Solve
(a) 〖 4〗^x=8,
4^x = 2^3
2^(2x) = 2^3
2x = 3
x = 3/2
===========================
(b) 〖 2〗^8÷8=2^(x-1)
2^8 + 8 = 2^(x-1)
--------
256+8 = 2^(x-1)
---------
264 = 2^(x-1)
---------
(x-1)*log(2) = log(264)
---------
x-1 = log(264)/log(2) = 8.0444
x = 9.0444
---------------------
Cheers,
Stan H.